# Differentially Private Bilevel Optimization

### v2

@@ 1 Introduction
Bilevel optimization is a fundamental framework for solving optimization objectives of hierarchical structure, in which constraints are defined themselves by an auxiliary optimization problem. Formally, it is defined as
@@
where $F:\mathbb{R}^{d_{x}}\to\mathbb{R}$ is referred to as the hyperobjective, $f:\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}\to\mathbb{R}$ as the upper-level (or outer) objective, and $g:\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}\to\mathbb{R}$ as the lower-level (or inner) objective. While bilevel optimization is well studied for over half a century (Bracken and McGill, 1973), it has recently received significant attention due to its diverse applications in machine learning (ML). These include hyperparameter tuning (Bengio, 2000; Maclaurin et al., 2015; Franceschi et al., 2017, 2018; Lorraine et al., 2020; Engstrom et al., 2025), meta-learning (Andrychowicz et al., 2016; Bertinetto et al., 2018; Rajeswaran et al., 2019; Ji et al., 2020), neural architecture search (Liu et al., 2018), invariant learning (Arjovsky et al., 2019; Jiang and Veitch, 2022), and data reweighting (Grangier et al., 2023; Fan et al., 2024; Pan et al., 2024). In these applications, both the upper and lower level objectives typically represent some loss over data, and are given by empirical risk minimization (ERM) problems over a dataset $S=\{\xi_{1},\dots,\xi_{n}\}\in\Xi^{n}:$111 It is possible for the datasets with respect to $f$ and $g$ to be distinct (e.g., validation/training), perhaps of different sizes. In this case, $S$ is the entire dataset, and $n$ is the total number of samples. Letting $f(\cdot;\xi_{i})=0$ or $g(\cdot;\xi_{i})=0$ for certain indices in order to exclude corresponding data points from either objective, will not affect our results.
$f(\mathbf{x},\mathbf{y}):=f_{S}(\mathbf{x},\mathbf{y})=\frac{1}{n}\sum_{i=1}^{n}f(\mathbf{x},\mathbf{y},\xi_{i})~{},\qquad g(\mathbf{x},\mathbf{y}):=g_{S}(\mathbf{x},\mathbf{y})=\frac{1}{n}\sum_{i=1}^{n}g(\mathbf{x},\mathbf{y},\xi_{i})~{}.$ (ERM)
Both of which are often empirical proxies of stochastic (i.e. population) losses with respect to a distribution $\mathcal{P}$ over $\Xi:$
@@
Nonetheless, private algorithms for bilevel optimization remain largely underexplored, and in particular, no gradient-based algorithm (aka first-order, which uses only gradient queries) that solves bilevel problems under DP, is known to date. This is no coincidence: until recently, no first-order methods with finite time guarantees were known even for non-private bilevel problems. This follows the fact (Ghadimi and Wang, 2018, Lemma 2.1) that under mild regularity assumptions, the so-called hypergradient takes the form:
@@
Consequently, directly applying a “gradient” method to $F$ requires inverting Hessians of the lower level problem at each time step, thus limiting applicability in contemporary high-dimensional applications. Following various approaches to tackle this challenge (see Section 1.2), recent breakthroughs were finally able to provide fully first-order methods for bilevel optimization with non-asymptotic guarantees (Liu et al., 2022; Kwon et al., 2023; Yang et al., 2023; Chen et al., 2024). These recent algorithmic advancements show promising empirical results in large scale applications, even up to the LLM scale of ${\sim}10^{9}$ parameters (Pan et al., 2024).
@@
The only prior method we are aware of for DP bilevel optimization was recently proposed by Chen and Wang (2024), which falls short in two main aspects. First, their algorithm only provides some privacy guarantee which cannot be controlled by the user. Moreover, it requires evaluating and inverting Hessians at each step, namely it is not first-order, which limits scalability; see Section 1.2 for further discussion.
@@ 1.1 Our contributions
We present DP algorithms that solve bilevel optimization problems whenever the the upper level is smooth but not necessarily convex, and the lower level problem is smooth and strongly-convex. These are the first algorithms to do so using only first-order (i.e. gradient) queries of the upper- and lower-level objectives, and moreover, are the first bilevel algorithms of any sort under standard central DP constraints. Our contributions can be summarized as follows:
@@
    Moreover, our algorithm is also applicable to the case where $\mathcal{X}\subsetneq\mathbb{R}^{d_{x}}$ is a non-trivial constraint set, which is common in applications of bilevel optimization. [^2] For instance, in data reweighting $\mathcal{X}$ is the probability simplex, and in hyperparameter tuning it is the hyperparameter space, which is typically constrained (e.g., non-negative reals). In the constrained setting, we obtain the same guarantee as above in terms of the projected hypergradient (see Section 2 for details).
@@
    $\left\|\nabla F_{\mathcal{P}}(\mathbf{x})\right\|=\widetilde{\mathcal{O}}\bigg{(}\Big{(}\frac{\sqrt{d_{x}}}{\epsilon n}\Big{)}^{1/2}+\Big{(}\frac{d_{x}}{n}\Big{)}^{1/2}+\Big{(}\frac{\sqrt{d_{y}}}{\epsilon n}\Big{)}^{1/3}\bigg{)}~{},$
@@ 1.2 Related work
Bilevel optimization was introduced by Bracken and McGill (1973), and grew into a vast body of work, with classical results focusing on asymptotic guarantees for certain specific problem structures (Anandalingam and White, 1990; Ishizuka and Aiyoshi, 1992; White and Anandalingam, 1993; Vicente et al., 1994; Zhu, 1995; Ye and Zhu, 1997). There exist multiple surveys and books covering various approaches for these problems (Vicente and Calamai, 1994; Dempe, 2002; Colson et al., 2007; Bard, 2013; Sinha et al., 2017).
@@
Ghadimi and Wang (2018) observed Eq. (1) under strong-convexity of the inner problem using the implicit function theorem, asserting that the hypergradient can be computed via inverse Hessians, which requires solving a linear system at each point. Many follow up works built upon this second-order approach with additional techniques such as variance reduction, momentum, Hessian sketches, projection-free updates, or incorporating external constraints (Amini and Yousefian, 2019; Yang et al., 2021; Khanduri et al., 2021; Guo et al., 2021; Ji et al., 2021; Chen et al., 2021; Akhtar et al., 2022; Chen et al., 2022; Tsaknakis et al., 2022; Hong et al., 2023; Jiang et al., 2023; Abolfazli et al., 2023; Merchav and Sabach, 2023; Xu and Zhu, 2023; Cao et al., 2024; Dagréou et al., 2024). Only recently, the groundbreaking result of Liu et al. (2022) proved finite-time convergence guarantees for a fully first-order method which is based on a penalty approach. This result was soon extended to stochastic objectives (Kwon et al., 2023), with the convergence rate later improved by (Yang et al., 2023; Chen et al., 2024), and also extended to constrained bilevel problems (Yao et al., 2024; Kornowski et al., 2024). The first-order penalty paradigm also shows promise for some bilevel problems in which the inner problem is not strongly-convex (Shen and Chen, 2023; Kwon et al., 2024; Lu and Mei, 2024), which is generally a highly challenging setting (Chen et al., 2024; Bolte et al., 2025). Moreover, Pan et al. (2024) provided an efficient implementation of this paradigm, showing its effectiveness for large scale applications.
@@
As to DP optimization, there is an extensive literature both for ERM and stochastic losses, which are either convex (Chaudhuri et al., 2011; Kifer et al., 2012; Bassily et al., 2014; Wang et al., 2017; Bassily et al., 2019; Feldman et al., 2020; Gopi et al., 2022; Carmon et al., 2023) or smooth and nonconvex (Wang et al., 2019; Tran and Cutkosky, 2022; Arora et al., 2023; Ganesh et al., 2024; Lowy et al., 2024).
@@
To the best of our knowledge, the only existing result for DP bilevel optimization is the recent result of Chen and Wang (2024), which differs than ours in several aspect. Their proposed algorithm is second-order, requiring evaluating Hessians and inverting them at each time step, which we avoid altogether. Moreover, Chen and Wang (2024) study the local DP model (Kasiviswanathan et al., 2011), in which each user (i.e. $\xi_{i}$) does not reveal its individual information. Due to this more challenging setting, they only derive guarantees for some finite privacy budget $\epsilon<\infty$, even as the dataset size grows. We study the common central DP model, in which a trusted curator acts on the collected data and releases a private solution, and we are able to provide any desired privacy and accuracy guarantees with sufficiently many samples. Our work is the first to study bilevel optimization in this common DP setting. We also note that Fioretto et al. (2021) studied the related problem of DP in Stackelberg games, which are certain bilevel programs which arise in game theory, aiming at designing coordination mechanisms that maintain each agent’s privacy.
@@ 2 Preliminaries
Two datasets $S,S^{\prime}\in\Xi^{n}$ are said to be neighboring, denoted by $S\sim S^{\prime}$, if they differ by only one data point. A randomized algorithm $\mathcal{A}:\Xi^{n}\to\mathcal{R}$ is called $(\epsilon,\delta)$ differentially private (or $(\epsilon,\delta)$-DP) for $\epsilon,\delta>0$ if for any two neighboring datasets $S\sim S^{\prime}$ and measurable $E\subseteq\mathcal{R}$ in the algorithm’s range, it holds that $\Pr[\mathcal{A}(S)\in E]\leq e^{\epsilon}\Pr[\mathcal{A}(S^{\prime})\in E]+\delta$ (Dwork et al., 2006). The basic composition property of DP states that the (possibly adaptive) composition of $(\epsilon_{0},\delta_{0})$-DP- and $(\epsilon_{1},\delta_{1})$-DP mechanisms, is $(\epsilon_{0}+\epsilon_{1},\delta_{0}+\delta_{1})$-DP. We next recall some well known DP basics: advanced composition, the Gaussian mechanism, and privacy amplification by subsampling.
Lemma 2.1 (Advanced composition, Dwork et al., 2010).
Lemma 2.2 (Gaussian mechanism).
Lemma 2.3 (Privacy amplification, Balle et al., 2018).
We remark that advanced composition will be used when $\epsilon_{0}\lesssim\sqrt{\log(1/\delta_{0})/T}$, thus the accumulated privacy scales as $\epsilon\asymp\sqrt{T}\epsilon_{0}$. Similarly, privacy amplification will be used when $\epsilon_{0}\leq 1$, under which the privacy after subsampling scales as $\epsilon\asymp\frac{b\epsilon_{0}}{n}$ (since $e^{\epsilon_{0}}-1\asymp\epsilon_{0},~{}(1-1/n)^{b}\asymp\frac{b}{n}$ and $\log(1+\frac{b}{n}\epsilon_{0})\asymp\frac{b}{n}\epsilon_{0}$).
@@
Given a point $\mathbf{x}\in\mathbb{R}^{d}$, and some $\mathbf{v}\in\mathbb{R}^{d},~{}\eta>0$, we denote
$\mathcal{G}_{\mathbf{v},\eta}(\mathbf{x}):=\frac{1}{\eta}\left(\mathbf{x}-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right)~{},~{}~{}~{}~{}~{}\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x}):=\arg\min_{\mathbf{u}\in\mathcal{X}}\left[\left\langle\mathbf{v},\mathbf{u}\right\rangle+\frac{1}{2\eta}\left\|\mathbf{u}-\mathbf{x}\right\|^{2}\right]~{}.$
In particular, given an $L$-smooth function $F:\mathbb{R}^{d}\to\mathbb{R}$ and $\eta\leq\frac{1}{2L}$, we denote the projected gradient (also known as reduced gradient) and the gradient (or prox) mapping, respectively, as
$\mathcal{G}_{F,\eta}(\mathbf{x}):=\frac{1}{\eta}\left(\mathbf{x}-\mathcal{P}_{\nabla F,\eta}(\mathbf{x})\right)~{},~{}~{}~{}~{}~{}\mathcal{P}_{\nabla F,\eta}(\mathbf{x}):=\arg\min_{\mathbf{u}\in\mathcal{X}}\left[\left\langle\nabla F(\mathbf{x}),\mathbf{u}\right\rangle+\frac{1}{2\eta}\left\|\mathbf{u}-\mathbf{x}\right\|^{2}\right]~{}.$
The projected gradient $\mathcal{G}_{F,\eta}(\mathbf{x})$ generalizes the gradient to the possibly constrained setting: for points $\mathbf{x}\in\mathcal{X}$ sufficiently far from the boundary of $\mathcal{X},~{}\mathcal{G}_{F,\eta}(\mathbf{x})=\nabla F(\mathbf{x})$, namely it simply reduces to the gradient; see the textbooks (Nesterov, 2013; Lan, 2020) for additional details. We will recall a useful fact, which asserts that the mapping $\mathcal{G}_{\mathbf{v},\eta}(\mathbf{x})$ is non-expansive with respect to $\mathbf{v}:$
Lemma 2.4.
For any $\mathbf{x},\mathbf{v},\mathbf{w}\in\mathbb{R}^{d},~{}\eta>0:~{}\left\|\mathcal{G}_{\mathbf{v},\eta}(\mathbf{x})-\mathcal{G}_{\mathbf{w},\eta}(\mathbf{x})\right\|\leq\left\|\mathbf{v}-\mathbf{u}\right\|$.
Lemma 2.4 is important in our analysis, since as we will argue later (in Section 4.1), gradient estimates must be inexact in the bilevel setting to satisfy privacy, and Lemma 2.4 will allows us to control the error due to this inexactness. Although this result is known (cf. Ghadimi et al. 2016), we reprove it in Appendix B for completeness.

### v3

@@ 1 Introduction
Bilevel optimization (BO) is a fundamental framework for solving optimization objectives of hierarchical structure, in which constraints are defined themselves by an auxiliary optimization problem. Formally, it is defined as
@@
where $F:\mathbb{R}^{d_{x}}\to\mathbb{R}$ is referred to as the hyperobjective, $f:\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}\to\mathbb{R}$ as the upper-level (or outer) objective, and $g:\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}\to\mathbb{R}$ as the lower-level (or inner) objective. While BO is well studied for over half a century (Bracken and McGill, 1973), it has recently received significant attention due to its diverse applications in machine learning (ML). These include hyperparameter tuning (Bengio, 2000; Maclaurin et al., 2015; Franceschi et al., 2017, 2018; Lorraine et al., 2020; Engstrom et al., 2025), meta-learning (Andrychowicz et al., 2016; Bertinetto et al., 2018; Rajeswaran et al., 2019; Ji et al., 2020), neural architecture search (Liu et al., 2018), invariant learning (Arjovsky et al., 2019; Jiang and Veitch, 2022), and data reweighting (Grangier et al., 2023; Fan et al., 2024; Pan et al., 2025). In these applications, both the upper and lower level objectives typically represent some loss over data, and are given by empirical risk minimization (ERM) problems over a dataset $S=\{\xi_{1},\dots,\xi_{n}\}\in\Xi^{n}:$ [^1] It is possible for the datasets with respect to $f$ and $g$ to be distinct (e.g., validation/training), perhaps of different sizes. In this case, $S$ is the entire dataset, and $n$ is the total number of samples. Letting $f(\cdot;\xi_{i})=0$ or $g(\cdot;\xi_{i})=0$ for certain indices in order to exclude corresponding data points from either objective, will not affect our results.
$f(\mathbf{x},\mathbf{y}):=f_{S}(\mathbf{x},\mathbf{y})=\frac{1}{n}\sum_{i=1}^{n}f(\mathbf{x},\mathbf{y},\xi_{i})~,\qquad g(\mathbf{x},\mathbf{y}):=g_{S}(\mathbf{x},\mathbf{y})=\frac{1}{n}\sum_{i=1}^{n}g(\mathbf{x},\mathbf{y},\xi_{i})~,$ (ERM)
where both objectives are often proxies of stochastic (i.e., population) losses with respect to a distribution $\mathcal{P}$ over $\Xi:$
@@
Nonetheless, private algorithms for BO remain largely underexplored, and in particular, no gradient-based algorithm (aka first-order, which uses only gradient queries) that solves BO problems under DP, is known to date. This is no coincidence: until recently, no first-order methods with finite time guarantees were known even for non-private bilevel problems. This follows the fact (Ghadimi and Wang, 2018, Lemma 2.1) that under mild regularity assumptions, the so-called hypergradient takes the form:
@@
Consequently, directly applying a “gradient” method to $F$ requires inverting Hessians of the lower-level problem at each time step, thus limiting applicability in contemporary high-dimensional applications. Following various approaches to tackle this challenge (see Section 1.2), recent breakthroughs were finally able to provide fully first-order methods for BO with non-asymptotic guarantees (Liu et al., 2022; Kwon et al., 2023; Yang et al., 2023; Chen et al., 2024). These recent algorithmic advancements show promising empirical results in large scale applications, even up to the LLM scale of ${\sim}10^{9}$ parameters (Pan et al., 2025).
@@
The only prior method we are aware of for DP BO was proposed by Chen and Wang (2024), which falls short in two main aspects. First, their algorithm only provides some privacy guarantee which cannot be controlled by the user. Moreover, it requires evaluating and inverting Hessians at each step, namely it is not first-order, which limits scalability; see Section 1.2 for further discussion.
@@ 1.1 Our contributions
We present DP algorithms that solve BO problems whenever the the upper-level is smooth but not necessarily convex, and the lower-level problem is smooth and strongly-convex. These are the first bilevel algorithms of any sort under standard central DP constraints, and in particular, are the first private algorithms to do so using only first-order (i.e., gradient) queries of the upper- and lower-level objectives. Our contributions can be summarized as follows:
@@
    Moreover, our algorithm is also applicable to the case where $\mathcal{X}\subsetneq\mathbb{R}^{d_{x}}$ is a non-trivial constraint set, which is common in applications of BO. [^2] For instance, in data reweighting $\mathcal{X}$ is the probability simplex, and in hyperparameter tuning it is the hyperparameter space, which is typically constrained (e.g., non-negative reals). In the constrained setting, we obtain the same guarantee as above in terms of the projected hypergradient (see Section 2 for details).
@@
    $\left\|\nabla F_{\mathcal{P}}(\mathbf{x})\right\|=\widetilde{\mathcal{O}}\bigg(\Big(\frac{\sqrt{d_{x}}}{\epsilon n}\Big)^{1/2}+\Big(\frac{\sqrt{d_{y}}}{\epsilon n}\Big)^{1/3}+\Big(\frac{d_{x}}{n}\Big)^{1/2}+\frac{1}{n^{1/4}}\bigg)~,$
@@ 1.2 Related work
BO was introduced by Bracken and McGill (1973), and grew into a vast body of work, with classical results focusing on asymptotic guarantees for certain specific problem structures (Anandalingam and White, 1990; Ishizuka and Aiyoshi, 1992; White and Anandalingam, 1993; Vicente et al., 1994; Zhu, 1995; Ye and Zhu, 1997). There exist multiple surveys and books covering various approaches for these problems (Vicente and Calamai, 1994; Dempe, 2002; Colson et al., 2007; Bard, 2013; Sinha et al., 2017).
@@
Ghadimi and Wang (2018) observed Eq. (1) under strong-convexity of the inner problem using the implicit function theorem, asserting that the hypergradient can be computed via inverse Hessians, which requires solving a linear system at each point. Many follow up works built upon this second-order approach with additional techniques such as variance reduction, momentum, Hessian sketches, projection-free updates, or incorporating external constraints (Amini and Yousefian, 2019; Yang et al., 2021; Khanduri et al., 2021; Guo et al., 2021; Ji et al., 2021; Chen et al., 2021; Akhtar et al., 2022; Chen et al., 2022; Tsaknakis et al., 2022; Hong et al., 2023; Jiang et al., 2023; Abolfazli et al., 2023; Merchav and Sabach, 2023; Xu and Zhu, 2023; Cao et al., 2024; Dagréou et al., 2024). Only recently, the groundbreaking result of Liu et al. (2022) proved finite-time convergence guarantees for a fully first-order method which is based on a penalty approach. This result was soon extended to stochastic objectives (Kwon et al., 2023), with the convergence rate later improved by (Yang et al., 2023; Chen et al., 2024, 2025), and also extended to constrained bilevel problems (Yao et al., 2024; Kornowski et al., 2024). The first-order penalty paradigm also shows promise for some bilevel problems in which the inner problem is not strongly-convex (Shen and Chen, 2023; Kwon et al., 2024; Lu and Mei, 2024), which is generally a highly challenging setting (Chen et al., 2024; Bolte et al., 2025). Moreover, Pan et al. (2025) provided an efficient implementation of this paradigm, showing its effectiveness for large scale applications.
@@
As to DP optimization, there is an extensive literature both for ERM and stochastic losses, for either convex objectives (Chaudhuri et al., 2011; Kifer et al., 2012; Bassily et al., 2014; Wang et al., 2017; Bassily et al., 2019; Feldman et al., 2020; Gopi et al., 2022; Carmon et al., 2023) as well as for nonconvex objectives (Wang et al., 2019; Tran and Cutkosky, 2022; Arora et al., 2023; Ganesh et al., 2024; Lowy et al., 2024; Zhang et al., 2024; Kornowski et al., 2025).
@@
To the best of our knowledge, the only existing result for DP BO prior to our work is the result of Chen and Wang (2024), which differs than ours in several aspect. Their proposed algorithm is second-order, requiring evaluating Hessians and inverting them at each time step, which we avoid altogether. Moreover, Chen and Wang (2024) study the local DP model (Kasiviswanathan et al., 2011), in which each user (i.e., $\xi_{i}$) does not reveal its individual information. Due to this more challenging setting, they only derive guarantees for some finite privacy budget $\epsilon<\infty$, even as the dataset size grows. We study the common central DP model, in which a trusted curator acts on the collected data and releases a private solution, and we are able to provide any desired privacy and accuracy guarantees with sufficiently many samples. Our work is the first to study BO in this common DP setting. We also note that Fioretto et al. (2021) studied the related problem of DP in Stackelberg games, which are certain bilevel programs which arise in game theory, aiming at designing coordination mechanisms that maintain each agent’s privacy.
After the initial version of this work became available on arXiv (Kornowski, 2024), Lowy and Liu (2025) further studied DP BO, proposing second-order algorithms which at the computational cost of inverting Hessians, are able to achieve gradient bounds which, interestingly, remove the dependence on the lower-level dimension $d_{y}$ presented in this work.
@@ 2 Preliminaries
Two datasets $S,S^{\prime}\in\Xi^{n}$ are said to be neighboring if they differ by only one data point. A randomized algorithm $\mathcal{A}:\Xi^{n}\to\mathcal{R}$ is called $(\epsilon,\delta)$ differentially private (or $(\epsilon,\delta)$-DP) for $\epsilon,\delta>0$ if for any two neighboring datasets $S\sim S^{\prime}$ and measurable $E\subseteq\mathcal{R}$ in the algorithm’s range: $\Pr[\mathcal{A}(S)\in E]\leq e^{\epsilon}\Pr[\mathcal{A}(S^{\prime})\in E]+\delta$ (Dwork et al., 2006). In Appendix A we recall some well known DP basics used in our analyses: composition and advanced composition, the Gaussian mechanism, and privacy amplification by subsampling.
@@
Given a point $\mathbf{x}\in\mathbb{R}^{d}$, and some $\mathbf{v}\in\mathbb{R}^{d},~\eta>0$, we denote $\mathcal{G}_{\mathbf{v},\eta}(\mathbf{x}):=\frac{1}{\eta}\left(\mathbf{x}-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right),~\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x}):=\arg\min_{\mathbf{u}\in\mathcal{X}}\big[\left\langle\mathbf{v},\mathbf{u}\right\rangle+\frac{1}{2\eta}\left\|\mathbf{u}-\mathbf{x}\right\|^{2}\big]$. In particular, given an $L$-smooth function $F:\mathbb{R}^{d}\to\mathbb{R}$ and $\eta\leq\frac{1}{2L}$, we denote the projected gradient (also known as reduced gradient) and the gradient (or proximal) mapping, respectively, as
$\mathcal{G}_{F,\eta}(\mathbf{x}):=\frac{1}{\eta}\left(\mathbf{x}-\mathcal{P}_{\nabla F,\eta}(\mathbf{x})\right)~,~~~~~\mathcal{P}_{\nabla F,\eta}(\mathbf{x}):=\arg\min_{\mathbf{u}\in\mathcal{X}}\Big[\left\langle\nabla F(\mathbf{x}),\mathbf{u}\right\rangle+\frac{1}{2\eta}\left\|\mathbf{u}-\mathbf{x}\right\|^{2}\Big]~.$
The projected gradient $\mathcal{G}_{F,\eta}(\mathbf{x})$ generalizes the gradient to the possibly constrained setting: for points $\mathbf{x}\in\mathcal{X}$ sufficiently far from the boundary of $\mathcal{X},~\mathcal{G}_{F,\eta}(\mathbf{x})=\nabla F(\mathbf{x})$, namely it simply reduces to the gradient; see the textbooks (Nesterov, 2013; Lan, 2020) for additional details.